Home‎ > ‎Crew 133‎ > ‎

108-GreenHabReport

XO Gartrelle to CapCom

The Greenhab is operating normally. Today's power outage (from 2:41PM to 3:53PM) in the Main Hab had no effect on Green hab operations. 

The fan is operating normally at 68 degrees. 

The onion, basil, cilantro seedlings and the Mars life shoot were given a small amount of water around 9AM.

The hydroponic seedlings were not given water today. The were placed in the green hab at 9AM returned to the upper Main hab for the night at 5PM. A photo of the hydroponic seedlings in the main hab from this evening is attached. 

No other plants were watered today. 

The heater was set to 40 degrees at 9AM
The heater was set to 75 degrees at 5PM

Greenhab readings for today are as follows:

                           Current                 Max                    Min   
Temperature    72 Degrees          88 Degrees         41 Degrees
Humidity              16%                      20%                    13%